# Diffraction of large-number whispering gallery mode by boundary straightening with jump of curvature

### v1

@@ 1 Introduction
$x=r\cos\varphi,\quad y=r\sin\varphi,\quad 0\leqslant r,\,\,-\pi\leqslant\varphi<\pi,$ (4)
$A$ is a diffraction coefficient, and $\mathcal{E}$ is a correction term which is not typically uniform in angle. Transition zones which surround limit rays where diffraction coefficients have singularities were not considered in detail.
@@
This work continues the research undertaken in [12]. Now, a much more complicated case is considered, when the number of transverse oscillations of the incident whispering gallery mode is large, but the parabolic equation method is still applicable. With its help, we study the wavefield in a small neighbourhood of the point of non-smoothness. Then, matching the obtained expressions with a cylindrical wave, we find the diffraction coefficient, see (2). It turns out that asymptotic formulas for the field in the case of incidence of the large-number whispering gallery mode are crucially different from the expressions obtained in the case of incidence of the small-number mode. As in the works [17, 18] devoted to a somewhat similar problem of diffraction at a inflection point, we pay a lot of attention to the ray structure of the field.
@@ 2 Problem statement
The wavefield $u$ is considered over the boundary $S=S_{-}\cup S_{+}$ which is an arc $S_{-}$ of circle of radius $a>0$, passing into a tangent straight line $S_{+}$ at the point $O$ (see Fig. 1), so that the curvature of $S$ jumps from $1/a$ to 00. There, $u$ satisfies the Helmholtz equation (1), and the Neumann condition holds on the boundary [^2] The case of the Dirichlet condition is treated similarly, and we will not dwell on it.
@@
We characterize the observation point $M=(x,y)$ positioned near $O$ by the distance $n\geqslant 0$ to the boundary along the normal and the arc length $s$ measured from $O$ (see Fig. 1), which is negative as $x<0$ and positive as $x>0$ (there, $s=x$ and $n=y$). Note that the mapping $(x,y)\to(s,n)$ is not smooth.
@@
Along the concave part of the boundary, an incident whispering gallery mode $u^{\text{inc}}$ runs to the point $O$. In a small neighbourhood of $O$, where
@@
$v(z)=$ $\frac{1}{2z^{1/4}}e^{-\frac{2}{3}z^{\frac{3}{2}}}\left(1+O\left(\frac{1}{z^{3/2}}\right)\right),\quad z\to+\infty,$ (10)
$v(z)=$ $\frac{1}{2i(-z)^{1/4}}e^{i\frac{2}{3}(-z)^{\frac{3}{2}}+i\frac{\pi}{4}}\left(1+O\left(\frac{1}{z^{3/2}}\right)\right)-$
-- $\frac{1}{2i(-z)^{1/4}}e^{-i\frac{2}{3}(-z)^{\frac{3}{2}}-i\frac{\pi}{4}}\left(1+O\left(\frac{1}{z^{3/2}}\right)\right),\quad z\to-\infty,$ (11)
@@
In this work, we aim at an asymptotic description of all waves arising to the right of the point of non-smoothness $O$ in its small neighbourhood. The research plan is as follows. First, a detailed geometrical analysis of the problem will be carried out, which allows to characterize areas with different wavefield structures. After that, we proceed to an analytical investigation. The wavefield in a small neighbourhood of the point $O$ will be studied within the framework of the parabolic equation method. The area of its applicability, as in the classical case [15, 16], is limited to the vicinity of the point $O$, where $s\ll 1$ and $n\ll 1$ (correspondingly, $x\ll 1$ and $y\ll 1$), but $\sigma$ and $\nu$ can be large. We will assume that inequalities
@@ 3.1 Qualitative discussion of ray structure of wavefield
of radius $a-b$ centred in the point $A$ (see Fig. 1). Bearing in mind limitation (14), we find
@@
The ray structure described above is retained up to the $y$ axis and collapses to the right of it.
@@
Before passing the caustic, the rays $\ell_{2}$ propagate above the ray $l_{O}$. After passing the caustic, they go to infinity between the rays $l_{B}$ and $l_{O}$. Thus, the ray $l_{O}$ (to the right of the point $Q$) is the boundary of light and shadow both for the waves $u_{1}$ and $u_{2}$. We call it the limit ray.
@@ 3.2 Rays of families $\ell_{1}$ and $\ell_{2}$
Let the observation point $M=(x,y)$ lie in a small neighbourhood of $O$. Consider a ray $l$ coming to $M$ immediately after its reflection from the concave part of the boundary $S_{-}$ (i.e. $l$ is either a ray of the family $\ell_{2}$ or a ray of the family $\ell_{1}$ before its reflection from the straight part of the boundary $S_{+}$). The case when $l$ comes to $M$ after reflection from $S_{+}$ will be considered later. We analytically describe position of the point $P=(0,h)$ of the intersection of $l$ with the $y$ axis and position of the point $T=(\xi,\eta)$ of the last reflection from $S_{-}$, see Fig. 3.
@@
1. Since the point $T$ is obviously close to $O$, its Cartesian coordinates $\xi$ and $\eta$ are expressed in terms of the corresponding arc length $s=\hat{s}$ (which is negative to the left of $O$):
@@
Note that $\alpha<0$ (respectively, $\hat{s}<-a\gamma$) for all the rays $\ell_{1}$ and $\alpha\geqslant 0$ (respectively, $\hat{s}\geqslant-a\gamma$) for all the rays $\ell_{2}$. For the ordinate of the point $P=(0,h)$, the equality
@@
Conditions $h\geqslant 0$ and $\hat{s}<0$ yield a limitation
$\hat{s}\geqslant-2a\gamma.$ (27)
Taking into account (24) and (23), an interesting relationship follows from (26):
@@
equals, in the main approximation, the distance from the observation point $M$ to the caustic $C$, see (16). $D$ is positive when $M$ lies below $C$, vanishes when $M$ is the caustic point, and is negative when $M$ lies above $C$ [^3] This agrees with the fact that the rays of the families $\ell_{1}$ and $\ell_{2}$ lie below the caustic. We are interested in negative values of $\hat{s}$. It is easy to show that for $x>a\gamma$ and $y\leqslant\gamma x$, i.e. for points $M$ lying to the right of $Q$ and below the limit ray $l_{O}$, see (20) and (21), equation (32) has one negative root
@@
becomes negative. In the case when $M$ does not lie in the areas specified above, equation (32) has no negative solutions. For $y<b$, i.e. for $M$ lying below $l_{B}$ (see Fig. 3), $\hat{s}_{1}<-a\gamma$ describes the position of the reflection point $T$ of the ray of the family $\ell_{1}$. If $y\geqslant b$, i.e. $M$ is located above $l_{B}$, then $\hat{s}_{1}\geqslant-a\gamma$ characterizes the point $T$ for the ray of the family $\ell_{2}$, which came to $M$ after passing the caustic. Expression (35) takes values $\hat{s}_{2}>-a\gamma$ and describes the position of the point $T$ if the ray of the family $\ell_{2}$ came to $M$ before passing caustic. Note that for $M$ lying on the caustic, $\widetilde{n}=0$ and $\hat{s}_{1}=\hat{s}_{2}$.
3. Position of the point $T$, at which the ray, arriving to the point $M$ after reflection from the straight part of the boundary, was last reflected from $S_{-}$, is determined quite similarly (the calculations differ from those above only by replacing $y$ with $-y$):
@@
characterizes the distance from the point $M^{\prime}$, symmetric to the observation point $M$ relative to $S_{+}$, to the caustic $C$. Condition (27) implies the inequality $y\leqslant\gamma x$, i.e. the point $M$ must lie below the limit ray $l_{O}$.
@@ 3.3 Eikonals of waves $u_{1}$ and $u_{2}$
To derive the formula for eikonal of $u_{1}$ after the rays $\ell_{1}$ reflect from $S_{+}$, it is enough to replace $\nu$ by $-\nu$ in (45).
@@
Note that inequalities (13) and (15) provide the smallness of correction terms in (45) and (46) (since $\nu-p=2\sigma\left(\sigma\pm\sqrt{\widetilde{\nu}}\right)$, see (38), (39)). Thus, it is clear that (13) and (15) can not be weaked.
@@ 4 Problem for parabolic equation and its exact solution
The main term of the asymptotics of the total wavefield is sought in a form similar to incident wave (9):
@@
Substituting (9) and (56) into the Helmholtz equation (1) and boundary condition (5), we obtain for $U$ the parabolic equation with a non-smooth coefficient
$iU_{\sigma}+U_{\nu\nu}+\nu\theta(-\sigma)U=0,$ (57)
@@
Here, $\theta$ is the Heaviside function, $\theta(z)=0$ as $z\leqslant 0$ and $\theta(z)=1$ as $z>0$.
An exact solution of (57), (58) is derived in [5, 12]. [^4] In [12], in contrast to [5], the detailed formulation of correct scattering problem for parabolic equation is given. It is unique in the class of functions which are quadratically integrable by $\nu$ and turn to the attenuation factor of the incident wave $U^{\text{inc}}$ (9) as $\sigma\to-\infty$. As $\sigma\leqslant 0$, it coincides with $U^{\text{inc}}$, and as $s>0$ it is given by the expression
[…]
@@ 5 Asymptotic investigation of attenuation factor
Since $\mathfrak{t}\gg 1$, the integration interval in (59) naturally splits out into three intervals where the intagrand behaves differently.
@@
Note that no splitting of the interval occurred in the case of $\mathfrak{t}=O(1)$ considered in [12].
@@ 1. We replace the Airy function with its asymptotics (11), and thus the integral $I_{1}$ becomes:
[…]
and the phases $f^{++}$, $f^{+-}$, $f^{-+}$ and $f^{--}$ of the exponentials are
@@
2. The asymptotics of each of the integrals $I_{1}^{\pm\pm}$ is given by contributions of the end point $p=0$ and the critical points of phases (63). The critical points of the phases are found from the equation
@@
1) the phase $f^{+-}(p)$ has one critical point $p_{1}=\mathfrak{t}-\left(\sigma-\sqrt{\widetilde{\nu}}\right)^{2}$ as $\mathfrak{t}>\nu$ (the quantity $\widetilde{\nu}$ is introduced in (41)).
@@
The expressions for the critical points $p_{1}$, $p_{2}$, $p_{3}$ coincide with expressions (38), (39), (40), respectively, for the points of intersection of the $y$ axis by rays of the families $\ell_{1}$ and $\ell_{2}$. The contributions of $p_{1}$, $p_{2}$, $p_{3}$ to the asymptotics of integrals $I^{\pm\pm}$, as we will see below, describe the contributions of the corresponding rays to the wavefield.
The exact formulation of the limitations on the the observation point coordinates $\sigma$ and $\nu$, ensuring that $p_{1}$, $p_{2}$, $p_{3}$ lie on the interval $L_{1}$ (60), is rather cumbersome, and we omit it. Note only that the inequalities $\mathfrak{t}-p_{1,2,3}\gg 1$ (due to the remarks above and relation (29)) can be interpreted as a condition that the rays corresponding to the critical points are not too close to the horizontal ray $l_{B}$, see Fig. 3.
3. Here are the limitations on the stretched coordinates $(\sigma,\nu)$ of observation points $M$ that guarantee the positivity of critical points, which are not difficult to obtain from the geometrical considerations carried out in section 3.2:
@@
2) $p_{1}\geqslant 0$ as $\sigma\leqslant\sqrt{\mathfrak{t}}$ and $\nu\geqslant 2\sqrt{\mathfrak{t}}\,\sigma$ (i.e., $M$ is positioned to the left of $Q$ and below the caustic, but above the limit ray $l_{O}$), and as $\sigma\geqslant\sqrt{\mathfrak{t}}$ and $\nu\leqslant 2\sqrt{\mathfrak{t}}\,\sigma$ (i.e., $M$ is positioned to the right of $Q$ and below the limit ray $l_{O}$).
3) $p_{2}\geqslant 0$ as $\sigma\leqslant\sqrt{\mathfrak{t}}$ and $\nu\geqslant 2\sqrt{\mathfrak{t}}\,\sigma$ (i.e., $M$ lies to the left of $Q$ and below the caustic, but above the limit ray $l_{O}$).
4) near the limit ray, where $\nu\approx 2\sqrt{\mathfrak{t}}\,\sigma$, the points $p_{1}$ and $p_{2}$ merge with the end point $p=0$: $p_{1}$ merges as $\sigma\geqslant\sqrt{\mathfrak{t}}$, and $p_{2}$ merges as $\sigma\leqslant\sqrt{\mathfrak{t}}$. If $M$ lies near caustic, where $\widetilde{\nu}$ (41) is small, then the points $p_{1}$ and $p_{2}$ merge. if $M$ is near the point $Q$ where caustic and the limit ray touch, then $p_{1}$, $p_{2}$ and $p=0$ merge all together.
5) $p_{3}\geqslant 0$ as $\nu\geqslant 2\sqrt{\mathfrak{t}}\,\sigma$ (i.e., $M$ is positioned below the limit ray $l_{O}$). As $\nu\approx 2\sqrt{\mathfrak{t}}\,\sigma$ (i.e., $M$ lies near the limit ray $l_{O}$) $p_{3}\approx 0$.

### v2

@@ 1 Introduction
$x=r\cos\varphi,\quad y=r\sin\varphi,\quad 0\leq r,\,\,-\pi\leq\varphi<\pi,$ (4)
AA is a diffraction coefficient, and $\mathcal{E}$ is a correction term which is not typically uniform in angle. Transition zones which surround limit rays where diffraction coefficients have singularities were not considered in detail.
@@
This work continues the research undertaken in [12]. Now, a much more complicated case is considered, when the number of transverse oscillations of the incident whispering gallery mode is large, but the parabolic equation method is still applicable. With its help, we study the wavefield in a small neighborhood of the point of non-smoothness. Then, matching the obtained expressions with a cylindrical wave, we find the diffraction coefficient, see (2). It turns out that asymptotic formulas for the field in the case of incidence of the large-number whispering gallery mode are crucially different from the expressions obtained in the case of incidence of the small-number mode. As in the works [17, 18] devoted to a somewhat similar problem of diffraction at a inflection point, we pay a lot of attention to the ray structure of the field.
@@ 2 Problem statement
The wavefield $u$ is considered over the boundary $S=S_{-}\cup S_{+}$ which is an arc $S_{-}$ of circle of radius $a>0$, passing into a tangent straight line $S_{+}$ at the point $O$ (see Fig. 1), so that the curvature of $S$ jumps from $1/a$ to 0. There, $u$ satisfies the Helmholtz equation (1), and the Neumann condition holds on the boundary [^2] The case of the Dirichlet condition is treated similarly, and we will not dwell on it.
@@
We characterize the observation point $M=(x,y)$ positioned near $O$ by the distance $n\geq 0$ to the boundary along the normal and the arc length $s$ measured from $O$ (see Fig. 1), which is negative as $x<0$ and positive as $x>0$ (there, $s=x$ and $n=y$). Note that the mapping $(x,y)\to(s,n)$ is not smooth.
@@
Along the concave part of the boundary, an incident whispering gallery mode $u^{\text{inc}}$ runs to the point $O$. In a small neighborhood of $O$, where
@@
$v(z)=$ $\frac{1}{2z^{\frac{1}{4}}}e^{-\frac{2}{3}z^{\frac{3}{2}}}\left(1+O\left(\frac{1}{z^{\frac{3}{2}}}\right)\right),\quad z\to+\infty,$ (10)
$v(z)=$ $\frac{1}{(-z)^{\frac{1}{4}}}\sin\left(\frac{2}{3}(-z)^{\frac{3}{2}}+i\frac{\pi}{4}\right)\left(1+O\left(\frac{1}{z^{3}}\right)\right)$ (11)
- $\frac{5}{24(-z)^{\frac{7}{4}}}\cos\left(\frac{2}{3}(-z)^{\frac{3}{2}}+i\frac{\pi}{4}\right)\left(1+O\left(\frac{1}{z^{3}}\right)\right)\quad z\to-\infty,$ (12)
@@
In this work, we aim at an asymptotic description of all waves arising to the right of the point of non-smoothness $O$ in its small neighborhood. The research plan is as follows. First, a detailed geometrical analysis of the problem will be carried out, which allows to characterize areas with different wavefield structures. After that, we proceed to an analytical investigation. The wavefield in a small neighborhood of the point $O$ will be studied within the framework of the parabolic equation method. The area of its applicability, as in the classical case [15, 16], is limited to the vicinity of the point $O$, where $s\ll 1$ and $n\ll 1$ (correspondingly, $x\ll 1$ and $y\ll 1$), but $\sigma$ and $\nu$ can be large. We will assume that inequalities
@@ 3.1 Qualitative discussion of ray structure of wavefield
of radius aba-b centred in the point $A$ (see Fig. 1). Bearing in mind limitation (15), we find
@@
The ray structure described above is kept up to the $y$ axis and collapses to the right of it.
@@
Before passing the caustic, the rays $\ell_{2}$ propagate above the ray $l_{O}$. After passing the caustic, they go to infinity between the rays $l_{B}$ and $l_{O}$. Thus, the ray $l_{O}$ (to the right of the point $Q$) is the shadow boundary both for the waves $u_{1}$ and $u_{2}$. We call it the limit ray.
@@ 3.2 Rays of families $\ell_{1}$ and $\ell_{2}$
Let the observation point $M=(x,y)$ lie in a small neighborhood of $O$. Consider a ray $l$ coming to $M$ immediately after its reflection from the concave part of the boundary $S_{-}$ (i.e. $l$ is either a ray of the family $\ell_{2}$ or a ray of the family $\ell_{1}$ before its reflection from the straight part of the boundary $S_{+}$). The case when $l$ comes to $M$ after reflection from $S_{+}$ will be considered later. We analytically describe position of the point $P=(0,h)$ of the intersection of $l$ with the $y$ axis and position of the point $T=(\xi,\eta)$ of the last reflection from $S_{-}$, see Fig. 3.
@@
1. Since the point $T$ is obviously close to $O$, its Cartesian coordinates $\xi$ and $\eta$ are expressed in terms of the corresponding arc length $\hat{s}$ (which is negative to the left of $O$):
@@
Note that $\alpha<0$ (respectively, $\hat{s}<-a\gamma$) for all the rays $\ell_{1}$ and $\alpha\geq 0$ (respectively, $\hat{s}\geq-a\gamma$) for all the rays $\ell_{2}$. For the ordinate of the point $P=(0,h)$, the equality
@@
Conditions $h\geq 0$ and $\hat{s}<0$ yield a limitation
$\hat{s}\geq-2a\gamma.$ (28)
Taking into account (24) and (25), an interesting relation follows from (27):
@@
equals, in the main approximation, the distance from the observation point $M$ to the caustic $C$, see (17). $D$ is positive when $M$ lies below $C$, vanishes when $M$ is the caustic point, and is negative when $M$ lies above CC. [^3] This agrees with the fact that the rays of the families $\ell_{1}$ and $\ell_{2}$ lie below the caustic. We are interested in negative values of $\hat{s}$. It is easy to show that for $x>a\gamma$ and $y\leq\gamma x$, i.e. for points $M$ lying to the right of $Q$ and below the limit ray $l_{O}$, see (21) and (22), equation (33) has one negative root
@@
becomes negative. In the case when $M$ does not lie in the areas specified above, equation (33) has no negative solutions. For $y<b$, i.e. for $M$ lying below $l_{B}$ (see Fig. 3), $\hat{s}_{1}<-a\gamma$ describes the position of the reflection point $T$ of the ray of the family $\ell_{1}$. If $y\geq b$, i.e. $M$ is located above $l_{B}$, then $\hat{s}_{1}\geq-a\gamma$ characterizes the point $T$ for the ray of the family $\ell_{2}$, which came to $M$ after passing the caustic. Expression (36) takes values $\hat{s}_{2}>-a\gamma$ and describes the position of the point $T$ if the ray of the family $\ell_{2}$ came to $M$ before passing caustic. Note that for $M$ lying on the caustic, $\widetilde{n}=0$ and $\hat{s}_{1}=\hat{s}_{2}$.
3. Position of the point $T$, at which the ray, arriving to the point $M$ after reflection from the straight part of the boundary, was last reflected from $S_{-}$, is determined quite similarly (the calculations differ from those above only by replacing $y$ with y-y):
@@
characterizes the distance from the point $M^{\prime}$, symmetric to the observation point $M$ relative to $S_{+}$, to the caustic $C$. Condition (28) implies the inequality $y\leq\gamma x$, i.e. the point $M$ must lie below the limit ray $l_{O}$.
@@ 3.3 Eikonals of waves $u_{1}$ and $u_{2}$
To derive the formula for eikonal of $u_{1}$ after the rays $\ell_{1}$ reflect from $S_{+}$, it is enough to replace $\nu$ by $\nu$ in (46).
@@
[…]
@@ 4 Problem for parabolic equation and its exact solution
The main term of the asymptotics of total wavefield is sought in a form similar to incident wave (9):
@@
Substituting (9) and (57) into the Helmholtz equation (1) and boundary condition (5), we obtain the parabolic equation with a non-smooth coefficient
$iU_{\sigma}+U_{\nu\nu}-\nu\theta(-\sigma)U=0,$ (58)
@@
Here, $\theta$ is the Heaviside function, $\theta(z)=0$ as $z\leq 0$ and $\theta(z)=1$ as $z>0$.
An exact solution of (58), (59) is derived in [5, 12]. [^4] In [12], in contrast to [5], the detailed formulation of correct scattering problem for parabolic equation is given. As $\sigma\leq 0$, it coincides with $U^{\text{inc}}$, and as $s>0$ it is given by the expression
$U=\frac{e^{-i\frac{\pi}{4}}}{2\sqrt{\pi\sigma}}\int_{0}^{\infty}v(p-\mathfrak{t})\left(e^{i\frac{(p-\nu)^{2}}{4\sigma}}+e^{i\frac{(p+\nu)^{2}}{4\sigma}}\right)dp.$ (60)
@@ 5 Asymptotic investigation of attenuation factor
Since $\mathfrak{t}\gg 1$, the integration interval in (60) naturally splits out into three intervals where the integrand behaves differently.
@@ 1. We replace the Airy function with its asymptotics (12), and thus the integral $I_{1}$ becomes:
$I_{1}^{-\pm}=\frac{1}{4\sqrt{\pi\sigma}}\int_{L_{1}}\frac{e^{if^{-\pm}(p)}}{(\mathfrak{t}-p)^{\frac{1}{4}}}\left(1+\frac{5i}{24(\mathfrak{t}-p)^{\frac{3}{2}}}+O\left(\frac{1}{(\mathfrak{t}-p)^{3}}\right)\right)dp,$
$I_{1}^{+\pm}=-\frac{i}{4\sqrt{\pi\sigma}}\int_{L_{1}}\frac{e^{if^{+\pm}(p)}}{(\mathfrak{t}-p)^{\frac{1}{4}}}\left(1-\frac{5i}{24(\mathfrak{t}-p)^{\frac{3}{2}}}+O\left(\frac{1}{(\mathfrak{t}-p)^{3}}\right)\right)dp,$
and the phases $^{++}, f+f^{+-}$, $f^{-+}$ and $f^{--}$ of the exponentials are
@@
2. The asymptotics of each of the integrals $I_{1}^{\pm\pm}$ is given by contributions of the end point $p=0$ and the critical points of phases (64).
The critical points of the phases are found from the equations
@@
1) the phase $f^{+-}(p)$ has one critical point $p_{1}=\mathfrak{t}-\left(\sigma-\sqrt{\widetilde{\nu}}\right)^{2}$ as $\mathfrak{t}>\nu$ ($\widetilde{\nu}$ is introduced in (42)).
@@
The expressions for the critical points $p_{1}$, $p_{2}$, $p_{3}$ coincide with expressions (39), (40), (41), respectively, for the points of intersection of the $y$ axis by rays of the families $\ell_{1}$ and $\ell_{2}$. The contributions of $p_{1}$, $p_{2}$, $p_{3}$ to the asymptotics of integrals $I_{1}^{\pm\pm}$, as we will see below, describe the contributions of the corresponding rays to the wavefield.
The exact formulation of the limitations on $\sigma$ and $\nu$, ensuring that $p_{1}$, $p_{2}$, $p_{3}$ lie on the interval $L_{1}$ (61), is rather cumbersome, and we omit it. Note only that the inequalities $\mathfrak{t}-p_{1,2,3}\gg 1$ (due to the remarks above and relation (30)) can be interpreted as a condition that the rays corresponding to the critical points are not too close to the horizontal ray $l_{B}$, see Fig. 3.
3. Here we list the limitations on the stretched coordinates $(\sigma,\nu)$ of observation points $M$ that guarantee the positivity of critical points, which are not difficult to obtain from the geometrical considerations carried out in section 3.2:
@@
2) $p_{1}\geq 0$ as $\sigma\leq\sqrt{\mathfrak{t}}$ and $\nu\geq 2\sqrt{\mathfrak{t}}\,\sigma$ (i.e., $M$ is positioned to the left of $Q$ and below the caustic, but above the limit ray $l_{O}$), and as $\sigma\geq\sqrt{\mathfrak{t}}$ and $\nu\leq 2\sqrt{\mathfrak{t}}\,\sigma$ (i.e., $M$ is positioned to the right of $Q$ and below the limit ray $l_{O}$).
3) $p_{2}\geq 0$ as $\sigma\leq\sqrt{\mathfrak{t}}$ and $\nu\geq 2\sqrt{\mathfrak{t}}\,\sigma$ (i.e., $M$ lies to the left of $Q$ and below the caustic, but above the limit ray $l_{O}$).
4) near the limit ray, where $\nu\approx 2\sqrt{\mathfrak{t}}\,\sigma$, the points $p_{1}$ and $p_{2}$ merge with the end point $p=0$: $p_{1}$ merges as $\sigma\geq\sqrt{\mathfrak{t}}$, and $p_{2}$ merges as $\sigma\leq\sqrt{\mathfrak{t}}$. If $M$ lies near caustic, where $\widetilde{\nu}$ (42) is small, then the points $p_{1}$ and $p_{2}$ merge. if $M$ is near the point $Q$ where caustic and the limit ray touch, then $p_{1}$, $p_{2}$ and $p=0$ merge all together.
5) $p_{3}\geq 0$ as $\nu\leq 2\sqrt{\mathfrak{t}}\,\sigma$ (i.e., $M$ is positioned below the limit ray $l_{O}$). As $\nu\approx 2\sqrt{\mathfrak{t}}\,\sigma$ (i.e., $M$ lies near the limit ray $l_{O}$) $p_{3}\approx 0$.